OpenStudy (anonymous):

When Raoul goes to the gym, he does 5 different exercises: bench press, squats, treadmill, leg press, and chin ups. If Raoul always does chin ups before he does bench press, how many different ways can he order his workout?

OpenStudy (kropot72):

The total number of ways of ordering the 5 different exercises must be divided by 2 the reason being that having fixed the positions of squats, treadmill and leg press, the chin ups and leg press can be arranged in only one way. Therefore the number of ways that the workout can be ordered is 5!/2.

OpenStudy (kropot72):

Can you calculate the value of (5 * 4 * 3 * 2 * 1)/2 ?

OpenStudy (kropot72):

\[\frac{5\times 4 \times 3 \times 2 \times 1}{2}=you\ can\ calculate\]

OpenStudy (anonymous):

That is a Brilliant.org Week question...